OpenStudy (anonymous):

how do i find the roots of x^6-1 and x^12-1and then plot them on a complex plane?

OpenStudy (anonymous):

sixth roots of one are ok. just find one of them. \[x^6=1\] \[x=1\] then divide the unit circle up into 6 equal parts

OpenStudy (anonymous):

a picture is the easiest way. you will get 1, -1, \[\frac{1}{2}+\frac{\sqrt{3}}{2}i\]

OpenStudy (anonymous):

the conjugate \[\frac{1}{2}-\frac{\sqrt{3}}{2}i\]

OpenStudy (anonymous):

also \[-\frac{1}{2}+\frac{\sqrt{3}}{2}i\]

OpenStudy (anonymous):

and its conjugate

OpenStudy (anonymous):

Okay, how did you get 1/2−3√/2i ? That seems to be where i am lost. Also, how do i split the unit circle into six compared to 4 or 8 sections of the unit circle

OpenStudy (anonymous):

ok are you working in degrees or radians? doesn't matter just let me k now

OpenStudy (anonymous):

u split a unit circle into 6 arcs of equal length

OpenStudy (anonymous):

We are doing both but I would like to first work with degrees. It is easier for me to undertstand

OpenStudy (anonymous):

see algebraically \[x^{6} = 1 = e^{i2k \pi}\]

OpenStudy (anonymous):

ok then you know that the circle has 360 degrees. divide this by 6 to get 60 degrees

OpenStudy (anonymous):

so \[x = e^{ik \pi/3}\]

OpenStudy (anonymous):

yeah satellites right

OpenStudy (anonymous):

so as you go around the circle you will go in multiples of 60 degrees:\0,60,120,180,240,300

OpenStudy (anonymous):

since you are on the unit circle the first coordinate will be cosine and the second coordinate will be sine

OpenStudy (anonymous):

cant we just do it algebraically and plot the points?

OpenStudy (anonymous):

i should be more accurate and say "the real part will be cosine and the imaginary part will be sine"

OpenStudy (anonymous):

@him yes of course but this is so easy if you have a picture of the unit circle in front of you. just divide into 6 parts and look at them

OpenStudy (anonymous):

bt then the algebra will give her a better idea of why were plotting the points

OpenStudy (anonymous):

I have the unit circle. How do i find cosine and sine

OpenStudy (anonymous):

see dizliz do u know how to plot e^(it)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

look at the circle on the last page. the first coordinate is cosine and the second coordinate is sine

OpenStudy (anonymous):

This is fairly straightforward in polar form....

OpenStudy (anonymous):

first try to understand whether she knows wht the polar form means and how to plot a complex no

OpenStudy (anonymous):

so, for example, at 60 degrees you see the point \[(\frac{1}{2},\frac{\sqrt{3}}{2})\]

OpenStudy (anonymous):

thinking of this as a point on the unit circle in the complex plane, the real part is \[\frac{1}{2}\] and the imaginary part is \[\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

\[as\ e^{i \pi/3} = \cos(\pi/3) + isin(\pi/3)\]

OpenStudy (anonymous):

so if you look at the cheat sheet i sent, you will see at all the multiples of 60 degrees the coordinates of the point. those are your answers. do you see where i mean?

OpenStudy (anonymous):

so the real part is cos (pi/3) and ther imaginary part is sin(pi/3)

OpenStudy (anonymous):

@him, no. @satellite I understand what you are saying about imaginary and real but not exactly sure where those numbers come from. Is there a formula?

OpenStudy (anonymous):

wht do u mean no? @dizliz @sat: i told u weve gotta explain

OpenStudy (anonymous):

see well show u do u agree \[1 = e^{i2k \pi}\]?

OpenStudy (anonymous):

where k is an integer

OpenStudy (anonymous):

are you asking where the numbers \[\frac{1}{2}\] amd \[\frac{\sqrt{3}}{2}\]come from?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lets explain satellite

OpenStudy (anonymous):

I know that a squared + b squared = y squared. Not sure if that helps any

OpenStudy (anonymous):

yes and i this case y = 1 since you are on the unit circle

OpenStudy (anonymous):

so you know that \[a^2+b^2=1\]

OpenStudy (anonymous):

yes that much I know

OpenStudy (anonymous):

as to where these particular coordinates come from, we can find them from scratch if you like. but i am presuming that before you were asked to find the 6th roots of 1 you had some trig yes?

OpenStudy (anonymous):

otherwise it is an unfair question. you should be familiar with basic trig before being asked to find roots of 1 in the complex plain

OpenStudy (anonymous):

i agree

OpenStudy (anonymous):

need to know the euler form and polar form

OpenStudy (anonymous):

Right now I am taking a summer course. I feel like all i do is take notes and he is s o unorganized so it is really hard to read his notes. I am taking precal 2 and this is the first I have ever seen any of this

OpenStudy (anonymous):

ok so lets start from scratch

OpenStudy (anonymous):

do we get into it satelite?

OpenStudy (anonymous):

he gave us euler's formula but with no explanation on how to use it or what it means

OpenStudy (anonymous):

there are certain basic numbers (angles) for which you should know the values of sine and cosine

OpenStudy (anonymous):

and in fact rather than derive these numbers which usually come from "reference triangles" let me ask you if by looking at the cheat sheet i sent, can you find \[\sin(120)\]?

OpenStudy (anonymous):

Assuming a knowledge of polar form (which is not hard to pick up if you have trig). Take z = 3 For z^3 = 1, there are 3 roots of unity. Represent z by <r, theta) and 1 by <1,0> so z^3 = <r^3, 3*theta> = <1,0> Taking the parts r^3 = 1 and 3*theta = 2m pi for some integer m. So z = < 1, 2m pi/3 > and you find the 3 roots with m = 1,2,3. To convert to Cartesian form <1, 2m pi/3> = cos ( 2m pi/3) + i sin ( 2m pi/3). This works generally not just for case = 3.

OpenStudy (anonymous):

i think we are starting with basic trig yes?

OpenStudy (anonymous):

were trying to explain the basics

OpenStudy (anonymous):

yes, it root of 3/2

OpenStudy (anonymous):

so now u know the trig ratios dizliz?

OpenStudy (anonymous):

ok and how did you find it?

OpenStudy (anonymous):

ratios?

OpenStudy (anonymous):

i meant from the cheat sheet, how did you find it?

OpenStudy (anonymous):

I think we must assume a knowledge of trig and the basics of complex numbers or this question is just inappropriate.

OpenStudy (anonymous):

Because it is the imaginary part that i seen at the 120 degree

OpenStudy (anonymous):

since shes only taking a summer course..we can excuse her

OpenStudy (anonymous):

second coordinate on the cheat sheet yes?

OpenStudy (anonymous):

yes second coordinate

OpenStudy (anonymous):

and how about \[\cos(300)\]?

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

ok and you did this just by looking at the unit circle. now your question may be "where did these numbers come from?" or it may be "how do i solve the problem i posted?" which one would you like to address first?

OpenStudy (anonymous):

ur amazing satellite

OpenStudy (anonymous):

where do these come from

OpenStudy (anonymous):

in other words without a cheat sheet how would i know they are there and what they are

OpenStudy (anonymous):

uve gotta memorize em

OpenStudy (anonymous):

ull just learn em over time

OpenStudy (anonymous):

ok that's good. but be warned that for practical purposes you should memorize something. let me find it for you

OpenStudy (anonymous):

do u get it dizliz>??

OpenStudy (anonymous):

dizliz?

OpenStudy (anonymous):

u there?>

OpenStudy (anonymous):

yes, I am here. So I just need to memorize the unit circle

OpenStudy (anonymous):

ok first of all do you know what a 30-60-90 triangle looks like?

OpenStudy (anonymous):

no you only have to memorize two triangles really

OpenStudy (anonymous):

Yes, that much I can gather

OpenStudy (anonymous):

a 30 -60 -90 triangle and a 45-45-90 triangle

OpenStudy (anonymous):

yes u need to memorize only some values

OpenStudy (anonymous):

not sure what you mean by 45-45-90

OpenStudy (anonymous):

lets do the easy one first. imagine you have a 45-45-90 triangle. since two sides are the same and it is a right triangle, if we set the hypotenuse = 1 (so it fits right inside the circle) you know that \[a^2+a^2=1\] \[2a^2=1\] \[a^2=\frac{1}{2}\] and so \[a=\sqrt{\frac{1}{2}}=\frac{1}{\sqrt{2}}=\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

means two angles are 45 degrees and it is a right triangle.

OpenStudy (anonymous):

now look at the coordinates of 45 degrees on the unit circle. you will see that both the first and second coordinate is \[\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

visualize placing the triangle in the circle so that the hypotenuse is the radius. those are the lengths of the sides

OpenStudy (anonymous):

Ok, I have a little bump where a goes from square root of 1/2 to square root of 2/2

OpenStudy (anonymous):

In radians, first figure is cos, second is sin

OpenStudy (anonymous):

that is the triangle on the upper right of the sheet i just sent

OpenStudy (anonymous):

oh let me unbump you

OpenStudy (anonymous):

\[\frac{1}{\sqrt{2}}=\frac{1}{\sqrt{2}}\times \frac{\sqrt{2}}{\sqrt{2}}=\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

this is called "rationalizing the denominator. they are the same numbers, just the form is different

OpenStudy (anonymous):

Okay, that's pretty easy.